# ConsistentEE: A Consistent and Hardness-Guided Early Exiting Method for Accelerating Language Models Inference

### 19 Dec 2023 v1

@@ Introduction
Why can camels survive for long without water?
Camels use the fat in their humps to keep them filled with energy and hydration for long periods of time.
Camels can survive for long without water because they can store water in their humps.
Camels can survival long periods without water due to their adaptation to desert conditions such as desertification of their habitat.
@@
Table 4: Snapshot of model responses under different speedup ratios.
Conclusion
We propose a reinforcement learning based approach to early exiting, so that at the training phase, only one internal classifier is required to predict the instance correctly. This makes the training phase consistent with the inference phase and can allow each layer to obtain better classification accuracy. For the reward function of the reinforcement learning framework, we propose the concept memorized layer to measure the hardness of each instance, and use it to dynamically balance accuracy and acceleration instead of using a fixed coefficient. Experimental results conducted on several commonly used datasets show that our approach is able to outperform the competitive baselines, demonstrating its effectiveness and efficiency.
Acknowledgement

### 07 Apr 2024 v2

@@ Acknowledgement
How should I prioritize my time??
Prioritize work that is tied to specific goals and projects.
Prioritize your time by investigating the needs of your daily life and focusing on the most crucial tasks.
Prioritize work that has a meaningful impact on your goals or vision.
Table 5: Snapshot of model responses under different speedup ratios on Alpaca/Dolly.
Context
Original Response
Response 2x
Response 5x
Five Americans who were monitored for three weeks at an Omaha, Nebraska, hospital after being exposed to Ebola in West Africa have been released, a Nebraska Medicine spokesman said in an email Wednesday. One of the five had a heart-related issue on Saturday and has been discharged but hasn’t left the area, Taylor Wilson wrote. The others have already gone home. They were exposed to Ebola in Sierra Leone in March, but none developed the deadly virus. They are clinicians for Partners in Health, a Boston-based aid group…
Five Americans who were exposed to Ebola in West Africa were released from a Nebraska hospital after three weeks of monitoring, with one still in the area for heart-related issue. The CDC expects to release the last of 17 patients by Thursday.
Five Americans exposed to Ebola in West Africa have been released from a Nebraska hospital after monitoring for three weeks. One has a heart-related issue and remains in the area, while the others have gone home.
Five Americans who were exposed to Ebola in Sierra Leone were released from a Nebraska hospital after three weeks of monitoring, with one still in the area for further evaluation.
(CNN)For the first time in eight years, a TV legend returned to doing what he does best. Contestants told to ”come on down!” on the April 1 edition of ”The Price Is Right” encountered not host Drew Carey but another familiar face in charge of the proceedings. Instead, there was Bob Barker, who hosted the TV game show for 35 years before stepping down in 2007…
Bob Barker returned to host ”The Price Is Right” for the first time in 8 years, filling in for Drew Carey. At 91, he looked spry and easily handled hosting duties, turning over to Carey at the end.
Bob Barker, former host of ”The Price is Right” returned to the show after 8 years, surprising contestants and hosting the first game with Drew Carey.
Bob Barker returned to hosting ”The Price is Right” for the first time in 8 years, filling in for Drew Carey during a taping of the show.
Table 6: Snapshot of model responses under different speedup ratios on CNN-DM.